OpenStudy (anonymous):

what is the least common mulitple of 7,8,10

OpenStudy (anonymous):

the easiest way is to make a chart like so:

OpenStudy (anonymous):

7: 14, 21, 28, 35, 42, 49, 56, etc 8: 16, 24, 32, 40, 48, etc 10: 20, 30, 40, 50, 60, etc and you do that untill you find the common multiple all three have

OpenStudy (anonymous):

it should be 280 , an faster way would be to use a factor tree i believe.

OpenStudy (anonymous):

7=7*1 8=2^3 10=5*2 then use all the result values like so: 7*2^3*5= 280